# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v1

@@ A.3 Finite Precision
In our negative results of Theorems 1 and 2 instead, we can pick the finite set $\mathbb{D}$ arbitrarily, e.g. floating point numbers, and we also make the use of the function $\mathrm{cast}:\mathbb{R}\to\mathbb{D}$, that we extend to $\mathbb{C}$ by applying it separately to real and imaginary part and to vector and matrices by applying it element-wise. The $\mathrm{cast}$ function is used because some computations of the state of the LRNN will be allowed to be in infinite precision and then transformed to finite precision using $\mathrm{cast}$ as specified in the proofs.
@@
We believe that the finite precision setup is not only realistic but also allows a better focus on the drawbacks of modern LRNN. Note that for Transformers, results usually rely instead on the notion of log-precision (Liu et al., 2023), meaning that the size of $\mathbb{D}$ grows logarithmically with the sequence length. This is mainly due to their limited expressivity compared to LRNNs. We also note that concerning the state-transition matrices of modern LRNNs (see Table 1), the values at the extremes of the eigenvalue range are technically not included (because of the use of the $\mathrm{sigmoid}$ and $\mathrm{softplus}$ functions). However, since we are working with finite precision, we can still include them by choosing the appropriate datatype $\mathbb{D}$, which in practice includes key values such as 00, $1$, and $-1$.
A.3.1 Initial State, Matrix-valued States, and The Decoder function
@@
When introducing the LRNN layer in (1), we mention that ${\bm{A}}$, ${\bm{B}}$ and $\mathrm{dec}$ are learnable functions. However, to learn the constructions in our theoretical results, we need also ${\bm{H}}_{0}\subseteq\mathbb{C}^{n\times d}$ to be learnable. We do this only to simplify the results since the same effect can also be achieved by using a special token \$currency-dollar\\$\$ at the beginning of each sequence input to the model, called the beginning of sequence token and setting, ${\bm{H}}_{0}=0$ for each LRNN layer so that ${\bm{B}}({\bm{x}}_{1})$ will have the same role as the learnable ${\bm{H}}_{0}$ in our constructions. This practice is standard and used in all our experiments.
@@
We also note that while we mention that the states ${\bm{H}}_{t}$ are generally matrices of dimension $n\times d$, for our theoretical constructions we always set $d=1$, so that states are vector-valued. Hence, for the problems that we consider, we find that having a matrix-valued state ($d>1$) brings no theoretical advantage.
@@
In (1), to compute the output $\hat{\bm{y}}_{t}$ from the state ${\bm{H}}_{t}$ and the vector ${\bm{x}}_{t}$ of an LRNN layer, we use the function $\mathrm{dec}$, to abstract away the computations that are done on ${\bm{H}}_{t}$ and ${\bm{x}}_{t}$, since they are not part of the recurrence. In this work, we do not consider the internal structure of $\mathrm{dec}$, but it usually contains a normalization and a feed-forward neural network and it can usually approximate any function.
@@
In our negative results on LRNNs expressivity in Theorems 1 and 2 our choice of arbitrary decoder guarantees the stronger results. For our positive results instead we either do not consider the decoder (Theorem 3) or we make use of a linear decoder (Theorem 4). We point out that to recognize regular languages efficiently and with a smaller LRNN state it is beneficial to have a more powerful (non-linear) decoder, as in the case of word problems for cyclic or permutation groups. However, such a decoder may be hard to approximate.
@@ Appendix B Parity and Modular Counting – Proofs
We report the full proofs for the theorems in Section 4.1.
@@ B.1 Proof of Theorem 1
The language $(11)^{*}$ contains all sequences with an even number of ones. An FSA recognizing the language, for the sequence $1^{k}$ will output $y_{k}=1$ if $k$ is even and $y_{k}=0$ if $k$ is odd. Consider an LRNN with one layer as in (1). We will prove that if the assumptions on the eigenvalues of ${\bm{A}}(1)$ are not satisfied, then there exists a $\overline{k}>0$ such that for every $k\geq\overline{k}$, the finite precision version of the state ${\bm{H}}_{k}$ corresponding to the sequence $1^{k}$ does not depend on $k$ and is equal to $\overline{{\bm{H}}}$. Hence, no matter the choice of $\mathrm{dec}$, also the finite precision version of $\hat{{\bm{y}}}_{k}$ will not vary with $k$ and thus for some $k^{\prime}\geq\bar{k}$, $\hat{\bm{y}}_{k^{\prime}}\neq k^{\prime}\mod 2=y_{k^{\prime}}$. An inductive argument can then be used for the case of LRNNs with multiple (finitely many) layers, using the fact that the input of the next layer will be constant for $k$ large enough, as the input of the first layers.
@@
where we set $\prod_{j=k}^{k-1}{\bm{A}}(x_{j})={\bm{I}}$ to avoid clutter. We follow Merrill et al. (2024) and make the simplifying assumption that in finite precision the state at time $k$ is computed by first evaluating all products involving the matrices ${\bm{A}}({\bm{x}}_{j})$ separately and in infinite precision, then casting them into finite precision, and finally executing the sum also in infinite precision and casting the result in finite precision. This avoids having to deal with matrix sums and products in finite precision. Hence, if we set ${\bm{x}}_{1}\dots\bm{x}_{k}=1^{k}$, we get the following exact and finite precision expressions for the state at time $k$.
@@
where cast is an operation that rounds matrices with complex values elementwise into finite precision by e.g. casting separately real and imaginary parts.
Using the Jordan canonical form theorem (see e.g. Horn & Johnson, 2012, Chap. 3.1) we can write ${\bm{A}}(1)={\bm{P}}{\bm{J}}{\bm{P}}^{-1}$, where ${\bm{J}}$ is block diagonal made of the Jordan blocks ${\bm{J}}_{1},\dots,{\bm{J}}_{s}$ with $s\leq n$, ${\bm{J}}_{i}\in\mathbb{R}^{k_{i}\times k_{i}}$ and with corresponding complex eigenvalues $\lambda_{1}\dots\lambda_{s}$ (with multiplicity taken into account). Such decomposition is useful because it allows to write matrix powers as
@@
Therefore, to study $\lim_{k\to\infty}{\bm{A}}(1)^{k}$, we can study the behavior of the elements of the Jordan blocks when $k\to\infty$. If $|\lambda_{i}|<1$ then all elements of ${\bm{J}}_{i}^{k}$ converge to zero, since the exponential is faster than the binomial $\binom{k}{j}$ with fixed $j$. Thus $\lim_{k\to\infty}{\bm{J}}_{i}^{k}=0$. If instead $\lambda_{i}\in\mathbb{R}$ and $\lambda_{i}>1$, then all nonzero elements of the Jordan block diverge to +$\infty$+ ∞. Finally, when $\lambda_{i}\in\mathbb{R}$ and $\lambda_{i}=1$, the diagonal elements are $\lambda_{i}^{k}=1$, while the other nonzero elements diverge to $\infty$. Therefore we have that if $|\lambda_{i}|<1$ or $\lambda_{i}$ is real and positive then there exists $\overline{\bm{J}}_{i}\in\{0,1,\infty\}^{k_{i}\times k_{i}}$ such that $\lim_{k\to\infty}{\bm{J}}_{i}^{k}=\overline{\bm{J}}_{i}$. Now, assume that for every $i$ either $|\lambda_{i}|<1$ or $\lambda_{i}\in\mathbb{R}$ with $\lambda_{i}\geq 1$. Then, from the structure of the Jordan decomposition, each element of the matrices ${\bm{A}}(1)^{k}{\bm{B}}(1)$ and ${\bm{A}}(1)^{k}{\bm{H}}_{0}$ will be a linear combination (with complex coefficients) of sequences of real numbers with well-defined limits (either $0,1$ or +$\infty$+ ∞), and thus, when $k\to\infty$ either converges to a point in $\mathbb{C}$ or diverges to a specific point in the complex infinity.
Now let $\widehat{\bm{C}}_{k}=\mathrm{cast}({\bm{A}}(1)^{k}{\bm{B}}(1))$ and $\widehat{\bm{D}}_{k}=\mathrm{cast}({\bm{A}}(1)^{k}{\bm{H}}_{0})$. Since $\mathrm{cast}$ operates elementwise and has a bounded and finite range we have that there exists $\tau\in\mathbb{N}$, $\widehat{\bm{C}}\in\mathbb{C}^{n\times d}$ and $\widehat{\bm{D}}\in\mathbb{C}^{n\times d}$ such that for every $k\geq\tau$, $\widehat{\bm{C}}_{k}=\widehat{\bm{C}}$ and $\widehat{\bm{D}}_{k}=\widehat{\bm{D}}$ and hence
@@
Note that only the matrix $k\widehat{\bm{C}}$ varies with $k$ and for $k\to\infty$ each element of $k\widehat{\bm{C}}$ has a well-defined limit, i.e. it converges to a point in $\overline{\mathbb{C}}$, that is the union of $\mathbb{C}$ and the complex infinity. It follows that each element of the matrix inside $\mathrm{cast}$ converges to a point in $\overline{\mathbb{C}}$. Therefore, since the $\mathrm{cast}$ operation has finite range we obtain that there exists $\overline{\bm{H}}\in\mathbb{C}^{n\times d}$ and $\bar{k}\geq\tau$ such that for every $k\geq\bar{k}$ we have $\widehat{\bm{H}}_{k}=\overline{\bm{H}}$, which concludes the proof. ∎

### v2

@@ A.4 Finite Precision
In our negative results of Theorems 1 and 2 instead, we can pick the finite set $\mathbb{D}\subset\mathbb{R}$ arbitrarily, e.g. floating point numbers, and we also make the use of the function $\mathrm{cast}:\mathbb{R}\to\mathbb{D}$, defined in (6). that we extend to $\mathbb{C}$ by applying it separately to the real and imaginary part and to vector and matrices by applying it element-wise. The $\mathrm{cast}$ function is used because some computations of the state of the LRNN will be allowed to be in infinite precision and then transformed to finite precision using $\mathrm{cast}$ as specified in the proofs. This function provides a simplification of the actual conversion that happens in practice.
@@
We believe that the finite precision setup is not only realistic but also allows a better focus on the drawbacks of modern LRNN. Note that for Transformers, results usually rely instead on the weaker notion of log-precision (Liu et al., 2023), meaning that the size of $\mathbb{D}$ grows logarithmically with the sequence length. This is mainly due to their limited expressivity compared to LRNNs. We also note that concerning the state-transition matrices of modern LRNNs (see Table 1), the values at the extremes of the eigenvalue range are technically not included (because of the use of the $\mathrm{sigmoid}$ and $\mathrm{softplus}$ functions). However, since we are working with finite precision, we can still include them by choosing the appropriate datatype $\mathbb{D}$, which in practice includes key values such as 00, $1$, and $-1$.
A.4.1 Initial State, Matrix-valued States, and The Decoder function
@@
When introducing the LRNN layer in (1), we mention that ${\bm{A}}$, ${\bm{B}}$ and $\mathrm{dec}$ are learnable functions. However, to learn the constructions in our theoretical results, we need also ${\bm{H}}_{0}\subseteq\mathbb{C}^{n\times d}$ to be learnable. We do this only to simplify the results, since the same effect can also be achieved by using a special token \$currency-dollar\\$\$ at the beginning of each sequence input to the model, called the beginning of sequence token and setting, ${\bm{H}}_{0}=0$ for each LRNN layer so that ${\bm{B}}({\bm{x}}_{1})$ will have the same role as the learnable ${\bm{H}}_{0}$ in our constructions. This practice is standard and used in all our experiments.
@@
While we mention that the states ${\bm{H}}_{t}$ are generally matrices of dimension $n\times d$, for our theoretical constructions (excluding the first two theorems), we set $d=1$, so that states are vector-valued. Hence, for the problems that we consider, we find that having a matrix-valued state ($d>1$) brings no theoretical advantage.
@@
To compute the output $\hat{\bm{y}}_{t}$ from the state ${\bm{H}}_{t}$ and the vector ${\bm{x}}_{t}$ of an LRNN layer in (1), we use the function $\mathrm{dec}$, to abstract away the computations that are done on ${\bm{H}}_{t}$ and ${\bm{x}}_{t}$, since they are not part of the recurrence. In this work, we do not consider the internal structure of $\mathrm{dec}$, but it usually contains a normalization and a feed-forward neural network and it can approximate any function.
@@
In our negative results on LRNNs expressivity in Theorems 1 and 2, our choice of an arbitrary decoder guarantees the stronger results. For our positive results instead, we either do not consider the decoder (Theorem 3) or we make use of a linear decoder (Theorem 4). We point out that to recognize regular languages efficiently and with a smaller LRNN state it is beneficial to have a more powerful (non-linear) decoder, as in the case of word problems for cyclic or permutation groups. However, such a decoder may be hard to learn.
@@ Appendix B Parity and Modular Counting – Proofs
We report the proofs for the theorems in Section 4.1. We start by defining the function $\mathrm{cast}:\mathbb{R}\to\mathbb{D}$, for a finite set $\mathbb{D}\subset\mathbb{R}$, which provides a simple model for the conversion of real numbers into a finite precision representation.
$\mathrm{cast}(x)=\min_{z\in\mathcal{D}_{\min}}z,\quad\mathcal{D}_{\min}:=\operatorname*{arg\,min}_{z\in\mathbb{D}}|z-x|.$ (6)
Note that $\mathcal{D}_{\min}$ might not be a singleton. We naturally extend this function on complex numbers by applying it separately to the real and imaginary part, and then to complex-valued matrices by applying it element-wise. The following lemma is a key element of the proofs of Theorems 1 and 2. The sequence $a_{k}$ in the lemma takes the form of the imaginary or real part of the elements of the $k$-th power of a matrix with real eigenvalues ($\lambda_{i}$ will be one eigenvalue), expressed using the Jordan canonical form. See Section B.1 for more details on the Jordan Canonical Form. Intuitively, the lemma shows that if some of the $\lambda_{i}$-s are negative then for $k$ large enough, $a_{k}$ in finite precision will alternate between two values. Instead, if the $\lambda_{i}$-s are only nonnegative, $a_{k}$ in finite precision becomes constant for large enough $k$.
Lemma 1.
Let $n,\bar{m}\in\mathbb{N}$ and for every $k>\bar{m}$ let
$a_{k}:=\sum_{i=1}^{n}c_{i}\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}},\quad\text{with }c_{i},\lambda_{i}\in\mathbb{R},m_{i}\in\mathbb{N},m_{i}\leq\bar{m},\quad\forall i\in\{1,\dots,n\},$
then there exist $\bar{k}\in\mathbb{N}$ such that for every $k\geq\bar{k}$ there exist $\bar{a}_{1},\bar{a}_{2}\in\mathbb{D}$ such that
$\mathrm{cast}(a_{2k})=\bar{a}_{1},\quad\mathrm{cast}(a_{2k+1})=\bar{a}_{2}.$
Furthermore, if $\lambda_{i}\geq 0$ for every $i\in\{1,\dots,n\}$, then $\mathrm{cast}(a_{k})=\bar{a}_{1}=\bar{a}_{2}$ for $k\geq\bar{k}$.
Proof.
If $c_{i}=0$ for every $i$, or $\lambda_{i}=0$ for every $i$, then $a_{k}=0$ for all $k$ and the statement is trivially satisfied. Without loss of generality we can assume that that $c_{i}\neq 0$ and $\lambda_{i}\neq 0$ for every $i\in\{1,\dots,n\}$, since for each $i$ where this is not true we can remove the corresponding term in the sum (since it will be 0) and use smaller value for $n$. We divide the proof into two parts.
Positive powers: Assume that $\lambda_{i}>0$ for all $i\in\{1,\dots,n\}$. This yields that for every $i$ and every $k{>}\bar{m}$, $\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}>0$. Since the $\mathrm{cast}$ function is piecewise constant with a finite number of pieces, we can divide the real line into a finite number of intervals where $\mathrm{cast}$ is constant. We now show that for $k$ large enough, the interval where $a_{k}$ belongs, and hence $\mathrm{cast}(a_{k})$, does not vary with $k$.
Without loss of generality we assume that for every $i,j\in\{1,\dots,n\}$ we have that $(m_{i},\lambda_{i})\neq(m_{j},\lambda_{j})$, since otherwise we can factor out $\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}$ and use a smaller $n$. Note that $\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}=\frac{k(k-1)\cdots(k-m_{i}+1)}{m_{i}!}\lambda_{i}^{k-m_{i}}$ and hence $g_{i}(k)=\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}$ for large $k$ behaves like the function $k^{m_{i}}\lambda^{k}$, i.e. the product of a polynomial and an exponential function of $k$. Without loss of generality, we therefore take the order of the indices of the terms in the sum such that the functions $g_{i}$ are in decreasing order of growth:
$\lambda_{i}>\lambda_{j}\text{ or }\lambda_{i}=\lambda_{j},m_{i}>m_{j}\qquad\forall i,j:i>j.$
By factoring out $g_{1}(k)$, i.e. the fastest growing term, from $a_{k}$ we get
$a_{k}=\binom{k}{m_{1}}\lambda_{1}^{k-m_{1}}\left(c_{1}+b_{k}\right)\qquad b_{k}:=\sum_{i=2}^{n}c_{i}\frac{\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}}{\binom{k}{m_{1}}\lambda_{1}^{k-m_{1}}},$
with $\lim_{k\to\infty}b_{k}=0$ and therefore, since for every $i$ and every $k>\bar{m}$, $\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}>0$ and $c_{1}\neq 0$, there exist $\hat{k}\in\mathbb{N}$ such that for every $k\geq\hat{k}$, $\operatorname{sign}(a_{k})=\operatorname{sign}(c_{1}+b_{k})=\operatorname{sign}(c_{1})$. Now let $\mathbb{D}=\{z_{1},\dots,z_{d}\}$ with $z_{1}<z_{2}<\dots<z_{d}$ and let $y_{1}=-\infty$, $y_{d+1}=\infty$ and $y_{i}=(z_{i-1}+z_{i})/2$ for $i\in\{2,\dots,d\}$. From its definition, $\mathrm{cast}$ is a piecewise constant function such that $\mathrm{cast}(x)=z_{i}$ for every $x\in(y_{i},y_{i+1})$. We now consider three cases according to the values of $\lambda_{1}$ and $m_{1}$.
1) If $\lambda_{1}>1$ or $\lambda_{1}=1,m_{1}>0$, then $\lim_{k\to\infty}\binom{k}{m_{1}}\lambda_{1}^{k-m_{i}}=\infty$ and there exists $\bar{k}\geq\hat{k}$ such that for every $k\geq\bar{k}$, either $a_{k}>y_{d}$ (if $\operatorname{sign}(c_{1})=1$) or $a_{k}<y_{2}$ (if $\operatorname{sign}(c_{1})=-1$) and hence $\mathrm{cast}(a_{k})=\bar{a}\in\{z_{1},z_{d}\}$.
2) If $\lambda_{1}<1$ then $\lim_{k\to\infty}\binom{k}{m_{1}}\lambda_{1}^{k-m_{i}}=0$ and hence there exist $\epsilon>0$, $j\in\{1,\dots,d\}$, $\bar{k}>\hat{k}$ such that for every $k\geq\bar{k}$, $a_{k}\in\Omega\subseteq(y_{j},y_{j+1})$, where $\Omega=(0,\epsilon)$ if $\operatorname{sign}(c_{1})=1$ and $\Omega=(-\epsilon,0)$ if $\operatorname{sign}(c_{1})=-1$. Therefore, $\mathrm{cast}(a_{k})=z_{j}$ for every $k\geq\bar{k}$.
3) If $\lambda_{1}=1,m_{1}=0$, then $\binom{k}{m_{1}}\lambda_{1}^{k-m_{i}}=1$ for every $k$ and hence
$a_{k}=c_{1}+b_{k},\quad b_{k}=\sum_{i=2}^{n}c_{i}\binom{k}{m_{i}}\lambda_{i}^{k-m_{i}}\qquad\text{with }\lambda_{i}<1\ \forall i\in\{2,\dots,n\}$
Note that $b_{k}$ has now the same structure as $a_{k}$, just with one less term in the sum, therefore we can factor out the term $\binom{\lambda_{2}}{m_{2}}\lambda^{k-m_{2}}$ and, since $\lambda_{2}<1$, apply the same reasoning as for the second case ($\lambda_{1}<1$) to $c_{1}+b_{k}$ and prove that there exist $\epsilon>0$, $j\in\{1,\dots,d\}$, $\bar{k}>\hat{k}$ such that for every $k\geq\bar{k}$, we have that $\operatorname{sign}(b_{k})=\operatorname{sign}(c_{2})$, $a_{k}\in\Omega\subseteq(y_{j},y_{j+1})$, where $\Omega=(c_{1},\epsilon)$ if $\operatorname{sign}(c_{2})=1$ and $\Omega=(-\epsilon,c_{1})$ if $\operatorname{sign}(c_{2})=-1$. Therefore $\mathrm{cast}(a_{k})=z_{j}$ for every $k\geq\bar{k}$.
In summary, we proved that when $\lambda_{i}\geq 0$ for every $i$, there exist $\bar{a}\in\mathbb{D}$, $\bar{k}\in\mathbb{N}$ such that for every $k\geq\bar{k}$ $a_{k}=\bar{a}$, which concludes the first part of the proof.
Some powers can be negative: Consider the general case where $\lambda_{i}\in\mathbb{R}$ can be negative. We can write
$a_{k}=\sum_{i=1}^{n}c_{i}\binom{k}{m_{i}}\mathrm{sign}(\lambda_{i})^{k-m_{i}}|\lambda_{i}|^{k-m_{i}}.$
Since $\mathrm{sign}(x)^{2k-m_{i}}$ and $\mathrm{sign}(x)^{2k+1-m_{i}}$ do not vary with $k$ we consider the two subsequences
$a_{2k}$ $=\sum_{i=1}^{n}\hat{c}_{i}\binom{2k}{m_{i}}|\lambda_{i}|^{2k-m_{i}},\quad\hat{c}_{i}=c_{i}\mathrm{sign}(\lambda_{i})^{2k-m_{i}}$
$a_{2k+1}$ $=\sum_{i=1}^{n}\tilde{c}_{i}\binom{2k+1}{m_{i}}|\lambda_{i}|^{2k+1-m_{i}},\quad\tilde{c}_{i}=c_{i}\mathrm{sign}(\lambda_{i})^{2k+1-m_{i}},$
and we can apply the same proof as for the case when $\lambda_{i}>0$ for every $i$ to each of the subsequences above, which gives the final result in the case $\lambda_{i}\in\mathbb{R}$ for every $i$. ∎
@@ B.1 Proof of Theorem 1
The language $(11)^{*}$ contains all sequences with an even number of ones. For the sequence of all ones $1^{k}$, an FSA recognizing the language, for the sequence $1^{k}$ will output $y_{k}=1$ if $k$ is even and $y_{k}=0$ if $k$ is odd. Consider an LRNN with one layer as in (1). We will prove that if ${\bm{A}}(1)$ has only nonnegative eigenvalues, then there exists a $\overline{k}>0$ such that for every $k\geq\overline{k}$, the finite precision version of the state ${\bm{H}}_{k}$ corresponding to the sequence $1^{k}$ does not depend on $k$ and is equal to $\overline{{\bm{H}}}$. Hence, no matter the choice of $\mathrm{dec}$, also the finite precision version of $\hat{{\bm{y}}}_{k}$ will not vary with $k$ and thus for some $k^{\prime}\geq\bar{k}$, $\hat{\bm{y}}_{k^{\prime}}\neq k^{\prime}\mod 2=y_{k^{\prime}}$. An inductive argument can then be used for the case of LRNNs with multiple (finitely many) layers, using the fact that the input of the next layer will be constant for $k$ large enough, as the input of the first layers.
@@
where we set $\prod_{j=k}^{k-1}{\bm{A}}(x_{j})={\bm{I}}$ to avoid clutter. We follow Merrill et al. (2024) and make the simplifying assumption that in finite precision the state at time $k$ is computed by first evaluating all products involving the matrices ${\bm{A}}({\bm{x}}_{j})$ separately and in infinite precision, followed by casting them into finite precision, and finally executing the sum also in infinite precision and casting the result in finite precision. This avoids having to deal with the individual matrix sums and products in finite precision, which would break associativity and be harder to analyze. Hence, if we set ${\bm{x}}_{1}\dots\bm{x}_{k}=1^{k}$, we get the following exact and finite precision expressions for the state at time $k$.
@@
where cast, defined in (6), is an operation that converts matrices with complex values element-wise into finite precision by e.g. separately converting real and imaginary parts.
Using the Jordan canonical form theorem (see e.g. Horn & Johnson, 2012, Chap. 3.1), we can write ${\bm{A}}(1)={\bm{P}}{\bm{J}}{\bm{P}}^{-1}$, where ${\bm{J}}$ is block diagonal made of the Jordan blocks ${\bm{J}}_{1},\dots,{\bm{J}}_{s}$ with $s\leq n$, ${\bm{J}}_{i}\in\mathbb{R}^{k_{i}\times k_{i}}$ and with corresponding complex eigenvalues $\lambda_{1}\dots\lambda_{s}$ (with multiplicity taken into account). Such decomposition is useful because it allows, for $k\geq\max_{i}k_{i}-1$, to write
@@
Then, from the structure of the Jordan decomposition, the imaginary and real part of each element of the matrices ${\bm{A}}(1)^{k}{\bm{B}}(1)$ and ${\bm{A}}(1)^{k}{\bm{H}}_{0}$ will be a linear combination of elements of the Jordan blocks taking the same form of $a_{k}$ in Lemma 1. Therefore since $\lambda_{i}\geq 0$ for every $i$, we can apply Lemma 1 component-wise and conclude that there exists $\tau\in\mathbb{N}$, $\widehat{\bm{C}}\in\mathbb{C}^{n\times d}$ and $\widehat{\bm{D}}\in\mathbb{C}^{n\times d}$ such that for every $k\geq\tau$, $\widehat{\bm{C}}_{k}=\mathrm{cast}({\bm{A}}(1)^{k}{\bm{B}}(1))=\widehat{\bm{C}}$ and $\widehat{\bm{D}}_{k}=\mathrm{cast}({\bm{A}}(1)^{k}{\bm{H}}_{0})=\widehat{\bm{D}}$ and hence
@@
Note that only the matrix $k\widehat{\bm{C}}$ varies with $k$ and for large enough $k$, the real and imaginary parts of each element of $k\widehat{\bm{C}}$ will be either 00, smaller than $\min_{x\in\mathbb{R}}\mathrm{cast}(x)$ or larger than $\max_{x\in\mathbb{R}}\mathrm{cast}(x)$. Therefore, we obtain that there exists $\overline{\bm{H}}\in\mathbb{C}^{n\times d}$ and $\bar{k}\geq\tau$ such that for every $k\geq\bar{k}$ we have $\widehat{\bm{H}}_{k}=\overline{\bm{H}}$, which concludes the proof. ∎